News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 28 July 2000 3.30pm EDT

During the last week, Chandra completed the observing schedule as planned.

The HETG was successfully retracted via stored commands, the first time this had been attempted since failure of the B-side retract limit switch on June 13. With the previous successful insertion, this now restores HETG usage to full normal operational capability.

The routine CAP 563A to check the CPE warm start count was run, count=1 as expected.

ASVT testing has progressed with certifying products needed for the manual LETG insertion/retraction operations, and for the lunar eclipse rehearsal of the solar array off-pointing.

The planned observations for the coming week are as follows:

------------------------------------------
Ophiuchi Mosaic #9      ACIS-I          Jul 31
SNR 0538-69.1           HRC-S
NGC 6397                ACIS-I
NGC 6500                ACIS-S          Aug 1
SN1998S                 ACIS-S
Comet c/LINEAR 1999     ACIS-S (16 pointings)
Radiation belts
LILLER 1                HRC-I           Aug 2
Elais:N2                ACIS-I
Elais:N1                ACIS-I          Aug 3
Radiation belts                         Aug 4
CL J2302.8+0844         ACIS-I          Aug 5
VEGA (3 calibrations)   HRC-S           Aug 6
G11.2-0.3               ACIS-S
GX 5-1                  ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports